OpenStudy (anonymous):

Suppose two angles, angle A and angle B, are adjacent complementary angles. Also assume that angle A is 30 degrees less than two times angle A and angle B. Find the measures of angle A and angle B

OpenStudy (lgbasallote):

angle A is 30 degrees less than two times angle A and angle B. ^what does that mean????

OpenStudy (lgbasallote):

30 less than twice the sum?

OpenStudy (anonymous):

* The second sentence is suppose to say, "Also assume that angle A is 30 degrees less than two times angle B"

OpenStudy (lgbasallote):

A + B = 90 A = 2B - 30 so 90 - B = 2b -30 120 = 3B B =40 A = 50 did you get the solution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks!